OpenStudy (kpop4life123):

. If y – 3 = 3x, which of the following sets represents possible inputs and outputs of the function, represented as ordered pairs? A {(0, 3), (1, 6), (2, 9)} B {(3, 0), (6, 1), (9, 2)} C {(1, 3), (2, 6), (3, 9)} D {(3, 1), (6, 2), (9, 3)

OpenStudy (math&ing001):

y - 3 = 3x is equivalent to y = 3x + 3. Let's plug in values of x, and check which of the ordered pairs represent the function. Example: if x=0, y=3*0+3=3 so the ordered pair (0,3) checks.

OpenStudy (kpop4life123):

how do we do it?

OpenStudy (math&ing001):

Just like I did with the ordered pair (0,3). Ordered pairs are written in the form of (x,y) What we do is we take the value of x and we plug it in the equation of y = 3x + 3. If the value of y that we find is equal to the one given, then the ordered pair represents the function.

OpenStudy (math&ing001):

All the pairs from a set must check so that we can say the set represents the function.